OpenStudy (anonymous):

Tell weather each equasion is true, false or open. 1. 62+(-15)=77 2. 52/-2=-16+(-30)/3 3. x/4+16=32-5/4

OpenStudy (anonymous):

1. 62 + (-15) = 77 62 - 15 = 77 47 = 77 (false) 2. 52/-2 = -16 + (-30)/3 -26 = -16 - 10 -26 = -26 (true) 3. x/4 + 16 = 32 - 5/4 open http://www.slideshare.net/gemmabean/introduction-to-equations